The One Product economy, which produces and sells only
personal computers (PCs), expects that it can sell 500 more, or 12,500
PCs, next year. Nominal GDP was $20 million this year, and the
money supply was $7 million. The central bank for the One Product
economy plans to increase the money supply by 10 percent next year.

a. What was the average selling price for the personal computers
this year?
b. What is the expected average selling price next year for
personal computers if the velocity of money remains at this
year’s turnover rate? What percentage change in price level is
expected to occur?
c. If the objective is to keep the price level the same next year
(i.e., no inflation), what percentage increase in the money
supply should the central bank plan for?
d. How would your answer in (c) change if the velocity of money
is expected to be three times next year? What is it now?
1
Expert's answer
2012-03-30T10:57:52-0400
a. P = GDP/Q = $20m/12,000 = $1,666.67
b. Velocity (V) was GDP/M = 20m/7m =
2.86, M will be 7m*1.1 = $7.7m, GDP = 7.7*2.86 = $22m, P = 22m/12,500 =
$1,760
%change = (1,760 - 1,666.67)/1,666.67 = 5.6%
c. M = P*Q/V =
1,666.67*12,500/2.86 = $7.28m, % = (7.28 - 7)/7 = 4%
d. M = P*Q/V =
1,666.67*12,500/(2.86*3) = $2.43m, % = (2.43 - 7)/7 = -65%